OpenStudy (anonymous):

The value of a computer in dollars, is given by c(t) =600 -150t, where t is the number of years since the computer was purchased. a. Find and interpret the vertical intercept of the fucntion c. b. Find and interpret the horizontal intercept of c. C. Find the domain and range of c.

OpenStudy (perl):

the vertical intercept, also known as y intercept, is the starting value of the computer at t=0 years

OpenStudy (unklerhaukus):

so sub in t=0 into c(t) c(t) = 600 -150t c(0) = 600 -150(0) and simplify

OpenStudy (unklerhaukus):

can you simplify this @cutegirl ? c(0) = 600 -150(0) = 600 -150*0 =

OpenStudy (anonymous):

600

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

can you interpret this "y intercept" in relation to the computer value?

OpenStudy (anonymous):

im kindof lost

OpenStudy (unklerhaukus):

think about what value the computer might be called when t=0

OpenStudy (anonymous):

t=600?

OpenStudy (unklerhaukus):

nope, think of some business type word

OpenStudy (unklerhaukus):

you got c(0)= 600 which is right

OpenStudy (anonymous):

ow ok i c

OpenStudy (anonymous):

then what

OpenStudy (unklerhaukus):

you have to interpret this in relation to the computer value equation,

OpenStudy (anonymous):

so 600 - 150?

OpenStudy (unklerhaukus):

that is c(1) = 600 - 150 = 450 which is the value one year after purchase

OpenStudy (anonymous):

thats for b?

OpenStudy (unklerhaukus):

not quite, the horizontal or x intercept is when the value of the computer = 0 ie c(t) = 600 -150t 0 = 600 -150t = = t =

OpenStudy (anonymous):

450

OpenStudy (unklerhaukus):

0 = 600 -150t 150t = 600 =

OpenStudy (anonymous):

450

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

no, divide 150t = 600 t = 600/150 t =

OpenStudy (anonymous):

t=4

OpenStudy (unklerhaukus):

yeah, the time that value of the computer reaches zero is 4 years